# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ 5.1 An extension with weaker notion of line connectivity
Let $\widetilde{\mathcal{H}}$ be the cell configuration constructed as above. The ergodicity modulo scaling property for $\widetilde{\mathcal{H}}$ holds by Lemma 5.4. Since each $\widetilde{H}_{H,k}$ contains the boundary of $H$, we see that the line connectivity property from Definition 1.4 holds. The almost planarity condition in Definition 1.5 also holds since one can add vertices and edges directly to the planar map that is close to $\mathcal{H}$ in the sense of Definition 1.5 and there are no macroscopic faces of $\mathcal{H}$ by Lemma $5.$ Furthermore, the moment bound (1.3) for $\widetilde{\mathcal{H}}$ follows from the moment bound (5.2). Therefore the conditions in Theorem 1.7 hold for $\widetilde{\mathcal{H}}$ and the conclusion follows from Theorem 1.7. ∎
5.3 Extension of Theorem 1.8 to $p$-angulations
In Theorem 1.8, we worked with the setting where the associated map $\mathcal{M}$ of the cell configuration $\mathcal{H}$ is a triangulation. In this section, we briefly explain how to extend Theorem 1.8 to the setting where $\mathcal{M}$ is a $p_{0}$-angulation for $p_{0}\geq 3$.
Let $\mathcal{M}$ be a one-ended planar map with no boundary. Assume that $\mathcal{M}$ has no loops, and every face of $\mathcal{M}$ is adjacent to $p_{0}$ vertices. We let $M(\mathcal{M})$ be the Riemann surface constructed in the same way as in the triangulation settings in Section 2.1, where we view each face of $\mathcal{M}$ as a regular $p_{0}$-gon with unit side length.
@@
Let $p_{0}\geq 3$. Let $\mathcal{H}$ be a random cell configuration such that its associated map is a one-ended $p_{0}$-angulation as above. Let $M(\mathcal{H})$ be the Riemann surface associated with the associated map of $\mathcal{H}$ defined as above. Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound (1.3). Then the conclusions in Theorem 1.8 still holds.
@@
The proof is almost identical to the proof of Theorem 1.8. The results in Section 2.2 still holds since we did not assume there that the associated map of $\mathcal{H}$ is a triangulation. To derive the regularity estimates in Section 4.1, one define the flowers and semi-flowers in the same way as in the triangulation case. It is straightforward to check that Lemma 4.2 still holds in this setting, and the rest of the proof in Section 4.1 still carries. To define the function $\phi_{0}$ in Section 4.2, we sample the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$ in the same way. For each regular $p_{0}$-gon on $M(\mathcal{H})$ formed by the cells $H_{1},...,H_{p_{0}}$, we further uniformly at random choose a way to draw straight lines to partition this face into $p_{0}-2$ triangles, and let $\phi_{0}$ be the piecewise linear function sending each such triangle on $M(\mathcal{H})$ to the corresponding triangle formed by the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$. Then one can check that for each face on $M(\mathcal{H})$ formed by $H_{1},...,H_{p_{0}}$, the Dirichlet energy of $\phi_{0}$ on this face is controlled by $C(p_{0})(\mathrm{diam}(H_{1})^{2}+....+\mathrm{diam}(H_{p_{0}})^{2})$ where $C(p_{0})$ is a constant depending only on $p_{0}$. This will allow us to bound the Dirichlet energy per area for $\phi_{0}$, and the rest of the proof is identical. ∎

### 06 May 2026 v2

@@ 5 Extensions to general planar maps
Let $\widetilde{\mathcal{H}}$ be the cell configuration constructed as above. The ergodicity modulo scaling property for $\widetilde{\mathcal{H}}$ holds by Lemma 5.2. Since each $\widetilde{H}_{H,k}$ contains the boundary of $H$, we see that the line connectivity property from Definition 1.4 holds. Furthermore, the moment bound (1.2) for $\widetilde{\mathcal{H}}$ follows from the moment bound (1.7). Therefore the conditions in Theorem 1.6 hold for $\widetilde{\mathcal{H}}$ and the conclusion follows from Theorem 1.6. ∎
5.2 Proof of Theorem 1.9
Let $\mathcal{M}$ be a planar map with whole-plane topology and let $\mathcal{H}$ be a cell configuration whose associated map is $\mathcal{M}$ as in Theorem 1.9. We assume $\mathcal{M}$ only has faces of degree at least 3; we may assume this upon removing loops and collapsing degree 2 faces as described above the theorem statement and observing that if the original map satisfies the assumptions of the theorem then the modified map with only face degrees at least 3 also satisfies these conditions. Let $\widetilde{\mathcal{T}}_{\mathcal{M}}$ be the triangulation generated by $\mathcal{M}$ as follows; see Figure 9 for an illustration. We add a vertex $v_{f}$ to each face $f\in\mathcal{F}\mathcal{M}$ and a vertex $v_{e}$ to each edge $e\in\mathcal{E}\mathcal{M}$. Adding the vertices $v_{e}$ cause each edge $e$ to be split into two edges, i.e., $e$ is replaced by two edges each connecting one end-point of $e$ and $v_{e}$. For each face $f$, we connect $v_{f}$ to all the vertices on $\mathcal{M}$ incident to the face $f$ and all the vertices $v_{e}$ where $e$ is an edge of $f$. Observe that $\widetilde{\mathcal{T}}_{\mathcal{M}}$ is an infinite plane triangulation without loops and multi-edges. Recall the construction of the cell configuration $\widetilde{\mathcal{H}}$ defined in terms of $\mathcal{H}$ as in Section 5.1. Let $d_{H}$ be the number of faces and edges of $\mathcal{M}$ incident to the vertex $v_{H}$. Let $(\widetilde{H}_{H,k})_{k=0,...,d_{H}}$ be the partition of the cell $H\in\mathcal{H}$ as defined in Section 5.1, except that we partition each cell into $d_{H}+1$ parts instead of $\deg(H)+1$ parts. For each edge $e$ (resp. face $f$) of $\mathcal{M}$ such that $v_{H}$ is on $e$ (resp. $f$), we uniformly randomly set $\widetilde{H}_{H,e}$ (resp. $\widetilde{H}_{H,f}$) to be equal to one of $\widetilde{H}_{H,k}$ for $k=1,...,d_{H}$. For each vertex $v_{f}$ (resp. $v_{e}$) added to generate $\mathcal{T}_{\mathcal{M}}$, let $\widetilde{H}_{f}=\cup_{H:v_{H}\sim f}\widetilde{H}_{H,f}$ (resp. $\widetilde{H}_{e}=\cup_{H:v_{H}\sim e}\widetilde{H}_{H,e}$), and let $\widetilde{H}_{H}=\widetilde{H}_{H,0}$. This defines a cell configuration $\widetilde{\mathcal{H}}$ with cells $\widetilde{H}_{f}$, $\widetilde{H}_{e}$ and $\widetilde{H}_{H}$ whose associated map is $\widetilde{\mathcal{T}}_{\mathcal{M}}$. We further define the conductance of edges $\widetilde{e}$ of $\widetilde{\mathcal{T}}_{\mathcal{M}}$ as follows. If $\widetilde{e}$ is connecting a vertex of $\mathcal{M}$ and a vertex $v_{e}$ in $\mathcal{V}\widetilde{\mathcal{T}}_{\mathcal{M}}$, we set $\mathfrak{c}(\widetilde{e})=1$. If $\widetilde{e}$ is connecting a vertex of $\mathcal{M}$ and a vertex $v_{f}$ in $\mathcal{V}\widetilde{\mathcal{T}}_{\mathcal{M}}$, we set $\mathfrak{c}(\widetilde{e})=\deg(f)+2$. If $\widetilde{e}$ is connecting two vertices $v_{e}$ and $v_{f}$, we set $\mathfrak{c}(\widetilde{e})=\deg(f)+1.5$. Note that by our choice of the conductance, the planar map $\mathcal{M}$ is measurable with respect to $(\widetilde{\mathcal{T}}_{\mathcal{M}},\mathfrak{c})$.
Recall the definition of semi-flowers in Section 4.1 in the setting where $\mathcal{M}$ is a triangulation. For a vertex $v$ of $\widetilde{\mathcal{T}}_{\mathcal{M}}$, we define the semi-flower $P_{v}\subseteq M(\mathcal{M})$ as follows. First suppose $v$ is also a vertex of $\mathcal{M}$, let $f$ be a face incident to $v$, and let $p\geq 3$ denote the degree of $f$. Let $Q_{f}$ be the unit side length regular $p$-gon associated with $f$, and we place $Q_{f}$ on the plane. Let $v_{1},...,v_{j_{v,f}}$ be vertices of $Q_{f}$ such that as we glue the polygon $Q_{f}$ according to the adjacency pattern of $\mathcal{M}$ as in the construction of $M(\mathcal{M})$, these vertices are identified with $v$. For each such vertex $v_{i}$, let $e_{1},e_{2}$ be the edges of $Q_{f}$ incident to $v_{i}$. We mark the points $x_{1},x_{2}$ on $e_{1},e_{2}$ such that they have distance $1/4$ from $v_{i}$. Let $e_{1}^{\perp},e_{2}^{\perp}$ be line segments started from $x_{1},x_{2}$ of length $\frac{1}{6\sqrt{3}}$ (the choice of this number is unimportant as long as it is less than $\frac{1}{4\sqrt{3}}$), perpendicular to $e_{1},e_{2}$, and pointing inwards to $Q_{f}$. Let $e_{12}$ be the line segment connecting the endpoints of $e_{1}^{\perp},e_{2}^{\perp}$ other than $x_{1},x_{2}$, and let $P_{v,f,i}\subseteq M(\mathcal{M})$ be the part of $Q_{f}$ enclosed by $e_{1},e_{2},e_{1}^{\perp},e_{2}^{\perp}$ and $e_{12}$ and we view $Q_{f}$ as part of $M(\mathcal{M})$. We let $P_{v}=\cup_{f\in\mathcal{F}\mathcal{M}:v\sim f}\cup_{i=1}^{j_{v,f}}P_{v,f,i}\subseteq M(\mathcal{M})$. Next suppose $v=v_{e}$ where $e$ is an edge on the face $f$ of $\mathcal{M}$. First assume $e$ is incident to $f$ and another face $\widetilde{f}$. Let $Q_{f}$ be the regular $p$-gon placed on the plane as before. Let $x_{3},x_{4}$ be points on $e$ of distance $1/4$ from the midpoint of $e$. Let $e_{3}^{\perp},e_{4}^{\perp}$ be line segments perpendicular to $e$ of length $\frac{1}{6\sqrt{3}}$, started from $x_{3},x_{4}$ and pointing inwards to $Q_{f}$. Let $e_{34}$ be the line segment connecting the endpoints of $e_{3}^{\perp},e_{4}^{\perp}$ other than $x_{3},x_{4}$. Let $P_{e,f}\subseteq M(\mathcal{M})$ be the corresponding part of enclosed by $e,e_{3}^{\perp},e_{4}^{\perp},e_{34}$ when we view $Q_{f}$ as part of $M(\mathcal{M})$. Define $P_{e,f^{\prime}}$ analogously, and let $P_{v_{e}}=P_{e,f}\cup P_{e,f^{\prime}}$. If $e$ is incident to the face $f$ only, then there is also some other edge $e^{\prime}$ of $Q_{f}$ identified with $e$ in our gluing of $Q_{f}$; in this case we define $P_{e,f}$ and $P_{e^{\prime},f}$ by drawing line segments on $Q_{f}$ in the same way, and let $P_{v_{e}}=P_{e,f}\cup P_{e^{\prime},f}$. Finally for a vertex $v_{f}$ where $f$ is a face of $\mathcal{M}$ of degree at least 3, let $P_{v_{f}}=Q_{f}\backslash\cup_{v\in\mathcal{V}\mathcal{M}:v\sim f}P_{v,f}$ be the remaining parts of $Q_{f}$. See Figure 9 for an illustration.
Refer to caption Refer to caption
Figure 9: Left: An illustration of part of $\widetilde{\mathcal{T}}_{\mathcal{M}}$, where the black edges are edges of the original map $\mathcal{M}$, and the brown edges are the edges added to construct $\widetilde{\mathcal{T}}_{\mathcal{M}}$. Right: An illustration of the semi-flowers on the surface $M(\mathcal{M})$. The green regions depict the semi-flowers $P_{v}$ where $v$ is a vertex of $\mathcal{M}$, the yellow regions depict the semi-flowers $P_{v_{f}}$ for $f\in\mathcal{F}\mathcal{M}$, and the blue regions are for the semi-flowers $P_{v_{e}}$ with $e\in\mathcal{E}\mathcal{M}$.
Lemma 5.4.
Let $M_{0}\subseteq M(\mathcal{M})$ be either equal to $M(\mathcal{M})$ or equal to the closure of a union of some faces of $\mathcal{T}_{\mathcal{M}}$ on $M(\mathcal{M})$ such that $M_{0}$ is homeomorphic to the closed unit disk. The conclusions of Lemma 4.1 hold for a conformal map $\varphi:M_{0}\to\mathbb{C}$ or $\varphi:M_{0}\to\mathbb{D}$ and the semi-flowers defined as above.
Proof.
Let $v$ be a vertex of $\mathcal{T}_{\mathcal{M}}$ on $M_{0}$. We only consider the setting where $v$ is in the interior of $M_{0}$; the boundary case is similar using Schwarz reflection. First suppose $v\in\mathcal{V}\mathcal{M}$. Let $r\in(0,\frac{{1}}{2})$. We define a cone $F_{v}(r)$ as follows. For each face $f$ incident to $v$ whose associated regular polygon is $Q_{f}$, let $F_{v,f}(r)$ be the intersection of $Q_{f}$ and the radius $r$ disks centered at the vertex $v_{i}$, where $v_{1},...,v_{j}$ are the vertices identified with $v$ in the gluing of $Q_{f}$. Let $F_{v}(r)=\cup_{f\in\mathcal{F}\mathcal{M}:v\sim f}F_{v,f}(r)$ and observe that $F_{v}(r)$ has the same geometry of a Euclidean cone of radius $r$ and some angle $\alpha_{v}\in[\frac{\pi}{3}\deg(v),\pi\deg(v))$, where we use that these disks are disjoint since $r<\frac{1}{2}$. Furthermore, from our construction $F_{v}(\frac{1}{16\sqrt{3}})\subseteq P_{v}\subseteq F_{v}(\frac{1}{{2\sqrt{3}}})\subseteq F_{v}(\frac{1}{3})\subseteq M(\mathcal{M})$. Therefore under the conformal map $\varphi_{v}:=z\mapsto(3{z})^{\frac{2\pi}{\alpha_{v}}}$, there are some universal constants $c_{0},c_{1}>0$ such that $B(0;c_{0})\subseteq\varphi_{v}(P_{v})\subseteq B(0;1-c_{1}\deg(v)^{-1})$, so the claim (i.e., (4.1)) follows from the Koebe distortion theorem. Now suppose $v=v_{f}$ where $f\in\mathcal{F}\mathcal{M}$ has degree $p\geq 3$. We place $Q_{f}$ on the plane such that $Q_{f}$ is centered at 0. Then
$B\Big(0;\cos(\frac{\pi}{p})(\frac{1}{2}\cot(\frac{\pi}{p})-\frac{1}{2\sqrt{3}})\Big)\subseteq P_{v_{f}}\subseteq B\Big(0;\sqrt{(\frac{1}{4})^{2}+(\frac{1}{2}\cot(\frac{\pi}{p})-\frac{1}{6\sqrt{3}})^{2}}\Big)\subseteq B(0;\frac{1}{2}\cot(\frac{\pi}{p}))\subseteq Q_{f},$
and in particular there are universal constants $c_{2},c_{3},c_{4}>0$ such that for some $r>0$, $B(0;c_{2}r)\subseteq P_{v_{f}}\subseteq B(0;c_{3}r(1-c_{4}p^{-1}))\subseteq B(0;c_{3}r)\subseteq Q_{f}$. The claim follows again from the Koebe distortion theorem. Finally, let $v=v_{e}$ for some edge $e\in\mathcal{E}\mathcal{M}$. Then from our construction of $M(\mathcal{M})$, on the surface $M(\mathcal{M})$, $P_{v_{e}}$ is a rectangle of side lengths $\frac{1}{2}$ and $\frac{1}{3\sqrt{3}}$ contained in the interior of a quadrangle of side length $\frac{1}{\sqrt{3}}$ and angles $\pi/3,2\pi/3,\pi/3,2\pi/3$, and the claim is again a consequence of the distortion theorem. ∎
Proof of Theorem 1.9.
Let $\widetilde{\mathcal{H}}$ be the cell configuration constructed as above. Then the moment bound (1.7) implies that in the cell configuration $\widetilde{\mathcal{H}}$,
$\mathbb{E}\big[\frac{\mathrm{diam}(\widetilde{H}_{0})^{2}}{\mathrm{area}(\widetilde{H}_{0})}\deg(\widetilde{H}_{0})^{4}\big]<\infty;\ \ \mathbb{E}\big[\frac{\mathrm{diam}(\widetilde{H}_{0})^{2}}{\mathrm{area}(\widetilde{H}_{0})}(\sum_{e:H_{0}\sim e}\mathfrak{c}(e))\big]<\infty.$ (5.5)
Furthermore, using identical arguments in Section 5.1, $\widetilde{\mathcal{H}}$ satisfies all the other conditions in Theorem 1.7. In particular, by the same arguments as in [GMS22, Lemma 2.10], for some deterministic constant $C>0$, for all large enough $k$,
$\sum_{H\in\widetilde{\mathcal{H}}(S_{k})}\mathrm{diam}(H)^{2}(\sum_{e:H_{0}\sim e}\mathfrak{c}(e))<C.$ (5.6)
Now the identical arguments in Section 4.1 for $\widetilde{\mathcal{H}}$ carry through once we use Lemma 5.4 instead of Lemma 4.1, and we define the function $\phi_{0}$ in Section 4.2 for $\widetilde{\mathcal{H}}$ in the identical way. It is elementary to prove that, for $H_{1},H_{2},H_{3}\in\widetilde{\mathcal{H}}$ a triangle $\Delta_{H_{1},H_{2},H_{3}}$ on $M(\mathcal{H})$ where $H_{3}$ is the cell for some vertex $v_{f}$, $\mathrm{Energy}(\phi_{0};\Delta_{H_{1},H_{2},H_{3}})\leq c_{0}\deg(H_{3})(\mathrm{diam}(H_{1})^{2}+\mathrm{diam}(H_{2})^{2}+\mathrm{diam}(H_{3})^{2})$ for some absolute constant $c_{0}>0$, which combined with (5.6) further implies the bound (4.34), i.e., the bound for the Dirichlet energy of $\phi_{0}$ as in Lemma 4.15. The rest of the proof is identical to that of Theorem 1.7. ∎
5.3 An extension with weaker notion of line connectivity
The following version of Theorems 1.6-1.9, which uses a weaker version of the line connectivity in Definition 1.4, will be used in our future work.
@@ 5.3 An extension with weaker notion of line connectivity
Suppose $\mathcal{H}$ is a cell configuration that satisfies all the conditions in one of the settings of Theorems 1.6-1.9 except for the line connectivity condition in Definition 1.4. Suppose there exists a cell configuration $\mathcal{H}^{\rm large}$ satisfying the condition in Definition 1.3 and the moment bound (1.2) along with the following properties:
  1. (i)
  2. (ii)
  3. (iii)
Then the conclusions in Theorem 1.6-1.9 continue to hold.
Theorem 5.5 can be viewed as a generalization of Theorem 1.6-1.9 since one can set $\mathcal{H}^{\rm large}=\mathcal{H}$ in the latter setting.
@@
[…]
The right side can be bounded exactly as [GMS22] since the arguments there only rely on the inputs from Lemma 2.3-2.7, which do not require the line connectivity property. Once this is done, the arguments in Section 2.3 and Section 3 work, except that in the construction of the map $\mathcal{M}_{k,\ell}$ in Lemma 2.14, we connect the cells $H_{j}$ and $H_{j+1}$ there using cells $H$ such that $\phi(H)$ intersects $L_{j,j+1}$.
In the setting of the Riemann uniformization, Definition 1.4 is used a few times. Let us explain the modifications. To prove Lemma 4.2 in this setting, we define
Then  (4.11) holds with $L^{\rm large}(\rho)$ instead of $L(\rho)$ using identical arguments applied to $\mathcal{H}^{\rm large}$, and (4.5) follows from identical arguments using conditions (ii) and (iii). The arguments for Lemma 4.24 are identical, except that at the final step we use that by Lemma 2.8 the set on the right hand side of (4.69) is contained within $\varphi(M(w+[-\delta 2^{k},\delta 2^{k}]^{2}))$, which is a connected subset of the set in (4.66). The rest of the proofs in Sections 4.1-4.4 do not require the line connectivity property and can be copied to the setting of Theorem 5.5. ∎